# Harmonic analysis of atomic vapor magnetization for in situ calibration of tri-axial time-dependent magnetic field

### v2

@@ Abstract
We introduce a methodology to calibrate in situ a set of coils generating bi- or tri-axial magnetic fields, at frequencies where a calibration performed under static conditions would be inaccurate. The coil constants are determined in a two-step procedure. Considering the presence of a static and a time-dependent field, firstly, the static one is oriented perpendicularly to the polarization plane of a time-dependent one; secondly, the polarization of the latter is made accurately circular. The methodology uses harmonic analysis of one component of the magnetization of an atomic sample whose spins adiabatically follow the time-dependent field.
preprint: preprint
I Introduction
@@
The works cited above consider time-dependent fields in the 100 Hz - 20 kHz range. It is worth noting that other research requires bi- or tri-axial fields at much higher frequencies, up to the radio [7] and microwave [8, 9, 10] ranges. We consider a methodology based on atomic spins that adiabatically follow the time-dependent field. This requirement limits the frequency at which the methodology can be practically applied to the audio range.
@@
The construction and characterization of magnetic field generators can follow both a priori and a posteriori approaches. The former involves designing the current distribution (coil shape and current waveform) aiming to produce the appropriate fields based on Maxwell’s laws. The a posteriori approach involves measuring the actual field and adjusting it using additional coils or varying the driving current(s). This can be essential for accuracy because the magnetic field generators are often surrounded by conducting materials or magnetic shields whose effects cannot be precisely considered in simulations or calculations. Several kinds of sensors can be used to measure the actual field, each of them coming with its inherent degree of precision, accuracy, spatial and time resolution, robustness etc. The use of Hall-effect or fluxgate sensors is often a favorite choice for their practicality, simple structure, and vector response [11, 12]. However, in some applications, they are not sufficiently precise or are not compatible with existing constraints.
Laser spectroscopy with its unrivaled precision is an area of research that can require extreme accuracy in magnetic field control. At the same time, it enables the construction of excellent magnetic detectors.
@@
As proposed by Breschi et al.[13], an elegant and effective way to control the field with the required accuracy and with detectors easily fitting in atomic spectroscopy setups is based on using the atomic sample itself as a magnetometric sensor. This concept forms the basis of in-situ calibration procedures reported by several researchers.
H.Zhang et al. [14] present a method to infer the coil constant (ratio between field and current) with high precision employing a high-performance hyperpolarized He magnetometer.
A similar calibration method is proposed by Yao et al. [15], who use a hybrid potassium-rubidium magnetometer and analyze the magnetic resonance under varying (swept) magnetic fields.
Chen et al. [16] propose a coil calibration method based on the duration of a $\pi/2$ pulse, which is precisely determined by the maximization of the initial amplitude of the free-induction-decay signal. The latter is generated by free precessing spins in Xenon gas polarized by collisions with optically pumped Rb vapour. This choice eliminates systematic errors that are possibly introduced by the laser radiation.
Zhao et al. [17] use calibrated coils to obtain vector response from a self-oscillating rubidium optically-pumped magnetometer non-destructively interrogated by far detuned probe radiation. This approach has similarities with that proposed by G.Zhang et al. [18], which uses a forced (Bell and Bloom) magnetometer. In this case, atoms are optically pumped by elliptically polarized radiation, which permits the achievement of a vector response from the analysis of orientation and alignment dynamics.
More recently K.Wang et al. [19] proposed a coil calibration method applied to a spin-exchange-relaxation-free optically-pumped magnetometer based on the dynamic (transient) response of such system upon application of sudden variation of the magnetic field along the three orthogonal directions.
A calibration method based on magnetic induction detection with an accurately aligned pickup coil has been proposed by K.Zhang et al. [20] for Helmholtz coils operated in a frequency range from few tens to several hundred Hz.
@@
The last assumption may be no longer valid in case of fast varying fields and currents. Spurious phenomena (e.g. parasitic capacitance of cables and coils) can cause the measured current across the series resistor to be different from that actually flowing in the coil, moreover, unperceived eddy currents may be induced in the coil proximity. Thus the high-frequency terms of the produced field may substantially differ, in amplitude and phase, from those inferred from the voltage drop measured on the monitor resistance [11, 1].
@@
The motivation that inspired the present work is the interest in studying the dynamics of atomic spins that evolve in (or are driven by) arbitrarily oriented time-dependent magnetic fields. In this research we have studied [21, 22] and applied [23] phenomena commonly referred to as magnetic dressing [24], focusing on cases where the dressing field contains various Fourier and Cartesian components [25, 6]. This research requires a very accurate control of the amplitude and relative phases of the magnetic field components. We present a method to accurately generate a circularly polarized field. This method permits an accurate a posteriori determination of the calibration parameters, which can subsequently be used to produce any kind of field polarization, over a broad frequency range, beyond the limit at which coil constants determined under static conditions become imprecise. The achieved results will find application in all cases where oscillating field components along diverse axes must be generated with very accurate relative phases and amplitudes.
@@
We consider a magnetic field ideally made of a time-dependent component that rotates on a plane and a static one oriented perpendicularly to that plane. Our goal is to detect imperfections of such field structure, which may consist of non-perfect perpendicularity of the static field and of elliptical polarization of the time-dependent one.
@@
The developed procedure is correspondingly made of two steps. The first makes the static field perpendicular to the polarization plane of the time-dependent one. The second one establishes a circular polarization of the latter.
@@
An optically-pumped atomic sample is interrogated employing polarimetric techniques, to produce a signal proportional to one component of the macroscopic magnetization. Diverse harmonic components of this signal bring information about the mentioned field imperfections, allowing an inherently in-situ calibration procedure to be developed based on Fourier analysis.
@@
The presentation is organized as follows: in Sec.II we describe the experimental setup, in Sec.III we describe (demanding calculus details to an Appendix) the principle of operation of the methodology to control the relative orientation of the static and the time-dependent fields and to refine the relative phases and amplitudes of the time-dependent field components. Demonstrative results are reported and analyzed in the Sec.IV. A synthesis of the achievements is drawn and shortly discussed in Sec.V.
II Setup
@@
The setup built to prepare and interrogate the atomic magnetization is described in Ref.[26]. Briefly (see Fig.1), Cs atoms are pumped employing circularly polarized radiation (at mW level) tuned to the D1 line and probed with a weak ($\mu$W level) linearly polarized radiation detuned by a few GHz from the D2 line. The D1 radiation is brought in resonance for tens of milliseconds and suddenly detuned to start the measurement. A balanced polarimeter is used to detect the Faraday rotation of the probe polarization plane, providing a signal proportional to the sample magnetization component along the probe-beam axis. The two beams co-propagate along the $x$ direction and in magnetometric application a static field is applied along the $z$ direction. The balanced polarimeter uses a transimpedance amplifier with a bandwidth of several kHz, enabling the detection of signals within the audio spectral range.
@@
Figure 1: Schematics of the setup. Cs atoms contained in a sealed cell (Cs) are optically pumped by circularly polarized radiation (D1) and probed by linearly polarized radiation (D2). The two laser beams co-propagate along $x$ and an interferential filter (IF) blocks the pump before the polarimeter (POL) that detects the Faraday rotation of the probe polarization. Three large-size Helmholtz pairs control the static magnetic field (only two of them, $B_{DCz}$ and $B_{DCx}$, are shown). Two smaller Helmholtz pairs ($B_{DCy}$ and $B_{DCz}$) generate time-dependent components along $y$ and $z$, while a solenoid ($B_{ACx}$) is used for the $x$. Several quadrupoles (not shown) help counteract static inhomogeneities.
@@
III Methodology
The proposed methodology requires that the atomic magnetization adiabatically follows the magnetic field to be characterized, i.e., the atomic spins must precess much faster than changes in field orientation. With a time-dependent field that rotates at an angular speed $\omega$, it is required that
$\gamma B\gg\omega,$ (1)
where $\gamma$ is the gyromagnetic factor and $B$ is the modulus of the field.
We consider a field nominally made of a static component along $z$ and a rotating one on the $xy$ plane. Possible imperfections are considered, which may concern the orientation of the DC field and some degree of ellipticity in the polarization of the time-dependent field. Namely, we consider
$\begin{split}\vec{B}&=B\hat{u}_{B}=\\ &=B_{0}\big{(}\cos(\omega t)+m_{1},(1+\varepsilon)\sin(\omega t+\phi)+m_{2},m_{3}\big{)}\end{split}$ (2)
with $m_{1},m_{2}$ accounting for spurious static field components on the polarization plane $xy$, and $\varepsilon,\phi$ for imperfections of the time-dependent field polarization. $B_{0}$ is the nominal amplitude of the rotating field and $m_{3}B_{0}$ is the nominal amplitude of the static one.
The aim is to obtain a time-dependent field with a circular polarization, i.e $\varepsilon=\phi=0$, on a plane perpendicular to the static term, i.e. $m_{1}=m_{2}=0$.
In the hypothesis of adiabatic following, the macroscopic magnetization $\vec{M}$ maintains a constant modulus $M$ (apart from the decay due to the relaxation mechanisms) while precessing around $\vec{B}(t)$. Among the components of $\vec{M}$, the one parallel to $\vec{B}$ remains approximately constant, while the perpendicular ones oscillate at the high frequency $\gamma|B|$ [30].
@@ III Methodology
The developed procedure is made of a first step aimed to make $m_{1}=m_{2}=0$, and a second one to make $\varepsilon=\phi=0$. Both the scopes can be pursued on the basis of Fourier analysis of the Taylor approximation of the quantity $M_{\parallel x}(t)$ (eq.(3)), the Appendix reports the relevant details.
III.1 Alignment of the static field along $z$ ($m_{1}=m_{2}=0$ condition)
Traditionally, the procedure to orient a static field along an axis (let it be $z$, as in our case) is based on minimizing the Larmor frequency by varying the currents that drive the $x$ and $y$ coils: the minimum is achieved when the $x$ and $y$ components of the field are fully compensated.
In the presence of a rotating field on the $xy$ plane, harmonic analysis of the detected signal (eq.(A.1)) can be used to reveal the presence of static components on the $xy$ plane, as well.
Let’s analyze the harmonic content of the measured signal described by eq.(3). The power ratio between the second-harmonic terms and the fundamental one (eqs.(A.3) and (A.2)) is:
@@ III.1 Alignment of the static field along $z$ ($m_{1}=m_{2}=0$ condition)
so that minimizing $A_{2}$ will lead to the condition $m_{1}=m_{2}=0$, i.e. to a static field perpendicular to the polarization plane of the time-dependent components.
Similar to the traditional procedure based on Larmor frequency minimization, the transverse components appear quadratically in the quantity $A_{2}$ to be minimized.
@@
As shown in the Appendix, non-zero values of $\phi$ and $\varepsilon$ do not contribute to second harmonic terms, making the described procedure robust to imperfect polarization of the rotating field. Indeed (see eqs. (A.3) and (A.4)), the polarization imperfections parametrized by $\phi$ and $\varepsilon$ produce only odd harmonic terms, as discussed in the next.
III.2 Refinement of the field circular polarization on the $xy$ plane ($\phi=\varepsilon=0$ condition)
Let’s assume that the static field (if any) has been aligned along $z$, i.e. let the condition $m_{1}=m_{2}=0$ and a generic $m_{3}$ be achieved. Now, harmonics of the fundamental tone are only ascribed to polarization imperfections, which, as shown in the Appendix, contribute only to odd harmonics of the detected signal.
Making reference to eqs.(A.2) and (A.4), we may derive the power ratio between the third harmonic terms and the fundamental one, finding
@@ III.2 Refinement of the field circular polarization on the $xy$ plane ($\phi=\varepsilon=0$ condition)
similarly to the case of eq.(4), minimizing $A_{3}$ will lead toward the $\varepsilon=\phi=0$ condition, i.e. to a condition of perfectly circular polarization of the time-dependent field.
It is worth examining the possible consequences of imperfect perpendicularity of the static field on the accuracy of this procedure. In fact, besides the analyzed second-harmonic terms, a non-perpendicular static field causes also third-harmonic signal, mimicking polarization issues. However, as shown in eq.(A.4), $m_{1}$ and $m_{2}$ in a first-order Taylor approximation do not produce such terms, and they only contribute when also the second order (eq.(A.5)) is taken into account. In this sense, the procedure aimed to obtain circular polarization is robust concerning the $m_{1}=m_{2}=0$ requirement.
IV Experimental validation
IV.1 Static field alignment
We have measured a FPS after having aligned the field along $z$ by Larmor frequency minimization and in the absence of any rotating field. The static field inferred from the FPS is about 1 $\mu$T. We have then applied a field of 16 $\mu$T rotating at 3144 Hz, thus the nominal condition is $m_{1}=m_{2}=0,\,m_{3}=1/16$, the amplitudes and the relative phase of the oscillating $x,y$ components are set to $\pi/2$ based on the static calibration factors. The recorded signal is reported in Fig.4: just a weak second harmonic peak appears, and third harmonic peak is visible, about 33 dB below the fundamental tone.

### v3

# Tri-axial time-dependent magnetic field calibrated in-situ by harmonic analysis of adiabatically evolving atomic spins

@@ Abstract
We introduce a methodology to calibrate in situ a set of coils generating bi- or tri-axial magnetic fields, at frequencies where a calibration performed under static conditions would be inaccurate. The methodology uses harmonic analysis of one component of the magnetization of an atomic sample whose spins adiabatically follow an ad hoc applied time-dependent field. The procedure enables the identification of phases and amplitudes of the coil currents required to produce a dynamic magnetic field with the assigned polarization. This determines coil constants that can be subsequently used to produce arbitrary three-dimensional time-dependent fields.
1 Introduction
@@
The characterization of magnetic field generators can follow both a priori and a posteriori approaches. The former involves designing the current distribution (coil shape and current waveform) aiming to produce the appropriate fields based on Maxwell’s laws. The a posteriori approach involves measuring the actual field and adjusting it using additional coils or varying the driving current(s). This can be essential for accuracy because the magnetic field generators are often surrounded by conducting materials or magnetic shields whose effects cannot be precisely considered in simulations or calculations. Several kinds of sensors can be used to measure the actual field, each of them coming with its inherent degree of precision, accuracy, spatial and time resolution, robustness etc. The use of Hall-effect or fluxgate sensors is often a favorite choice for their practicality, simple structure, and vector response [7, 8]. However, in some applications, they are not sufficiently precise, do not possess the necessary bandwidth, or are not compatible with existing constraints. In this respect, the methodology proposed in this work might help build setups for the calibration of those vectorial sensors.
@@
Laser spectroscopy with its unrivaled precision is an area of research that can require extreme accuracy in magnetic field control. At the same time, it enables the construction of excellent magnetic detectors. As proposed by Breschi et al.[9], an elegant and effective way to control the field with the required accuracy and with detectors easily fitting in atomic spectroscopy setups is based on using the atomic sample itself as a magnetometric sensor. This concept forms the basis of in-situ calibration procedures reported by several researchers.
H.Zhang et al. [10] presented a method to infer the coil constant (ratio between field and current) with high precision employing a high-performance hyperpolarized He magnetometer. A similar calibration method was proposed by Yao et al. [11], who used a hybrid potassium-rubidium magnetometer and analyze the magnetic resonance under varying (swept) magnetic fields. Chen et al. [12] proposed a coil calibration method based on the duration of a $\pi/2$ pulse, precisely determined by the maximization of the initial amplitude of the free-induction-decay signal. The latter was generated by free precessing spins in Xenon gas polarized by collisions with optically pumped Rb vapour. This choice eliminates systematic errors that are possibly introduced by the laser radiation. Zhao et al. [13] used calibrated coils to obtain vector response from a self-oscillating rubidium optically-pumped magnetometer non-destructively interrogated by far-detuned probe radiation. This approach has similarities with that proposed by G.Zhang et al. [14], which used a forced (Bell and Bloom) magnetometer. In this case, atoms were optically pumped by elliptically polarized radiation, enabling of a vector response from the analysis of orientation and alignment dynamics. More recently, K.Wang et al. [15] proposed a coil calibration method applied to a spin-exchange-relaxation-free optically-pumped magnetometer based on the dynamic (transient) response of such system upon application of sudden variation of the magnetic field along the three orthogonal directions. A calibration method based on magnetic induction detection with an accurately aligned pickup coil has been proposed by K.Zhang et al. [16] for Helmholtz coils operated in a frequency range from few tens to several hundred Hz.
@@
The last assumption may be no longer valid in the case of fast-varying fields and currents. Spurious phenomena (e.g. parasitic capacitance of cables and coils) can cause the measured current across the series resistor to be different from that actually flowing in the coil, moreover, unperceived eddy currents may be induced in the coil proximity. Thus the high-frequency terms of the produced field may substantially differ, in amplitude and phase, from those inferred from the voltage drop measured on the monitor resistance [7, 1].
@@
The motivation that inspired the present work is the interest in studying the dynamics of atomic spins that evolve in (or are driven by) arbitrarily oriented time-dependent magnetic fields. In this research we have studied [17, 18] and applied [19] phenomena commonly referred to as magnetic dressing [20], focusing on cases where the dressing field contains various Fourier and Cartesian components [21, 6]. This research requires a very accurate control of the amplitude and relative phases of the magnetic field components.
Here, we present an original method to accurately generate a circularly polarized field. This method permits an accurate a posteriori determination of the calibration parameters, which can subsequently be used to produce any kind of field polarization, over a broad frequency range, beyond the limit at which coil constants determined under static conditions become imprecise. Our method relies on atomic spins that follow adiabatically a time-dependent field and this requirement sets an upper limit to the practically applicable frequencies, which must approximately fall in the audio range. The achieved results will find application in all cases where oscillating field components along diverse axes must be generated with very accurate relative phases and amplitudes.
@@
The calibration methodology is based on generating a magnetic field ideally made of a time-dependent component that rotates on a plane and a static one oriented perpendicularly to that plane. This field drives the evolution of atomic spins and its imperfections, which may consist of non-perfect perpendicularity of the static field and of elliptical polarization of the time-dependent one, are pointed out by harmonic analysis of the measured signal.
@@
The developed procedure is correspondingly made of two steps. A preliminary step makes the static field perpendicular to the polarization plane of the time-dependent one. The second one establishes a circular polarization of the latter.
@@
An optically-pumped atomic sample is interrogated with polarimetric techniques extracting a signal proportional to one component of the macroscopic magnetization. Diverse harmonic components of this signal bring information about the mentioned field imperfections, allowing an inherently in-situ calibration procedure to be developed based on Fourier analysis.
@@
The presentation is organized as follows: in Sec.2 we describe the experimental setup, in Sec.3 we describe (demanding calculus details to an Appendix) the principle of operation of the methodology to control the relative orientation of the static and the time-dependent fields (Sec.3.1) and to refine the relative phases and amplitudes of the time-dependent field components (Sec.3.2). Demonstrative results are reported and analyzed in the Sec.4. A synthesis of the achievements is drawn and shortly discussed in Sec.5.
2 Setup
@@
The setup built to prepare and interrogate the atomic magnetization is described in Ref.[22]. Briefly (see Fig.1), Cs atoms are optically pumped employing circularly polarized radiation (at mW level) tuned to the D1 line, and probed with a weak ($\mu$W level) linearly polarized radiation detuned by a few GHz from the D2 line. The D1 radiation is brought in resonance for tens of milliseconds and suddenly detuned to start the measurement. A balanced polarimeter is used to detect the Faraday rotation of the probe polarization plane, providing a signal proportional to the sample magnetization component along the probe-beam axis. The two beams co-propagate along the $x$ direction and in magnetometric application a static field is applied along the $z$ direction. The balanced polarimeter uses a transimpedance amplifier with a bandwidth of several kHz, enabling the detection of signals within the audio spectral range.
@@
Figure 1: Schematics of the setup. Cs atoms contained in a sealed cell (Cs) are optically pumped by circularly polarized radiation (D1) and probed by linearly polarized radiation (D2). The two laser beams co-propagate along $x$ and an interferential filter (IF) blocks the pump before the polarimeter (POL) that detects the Faraday rotation of the probe polarization. Three large-size Helmholtz pairs control the static magnetic field (only two of them, $B_{\mathrm{DC}z}$ and $B_{\mathrm{DC}x}$, are shown). Two smaller Helmholtz pairs ($B_{\mathrm{AC}y}$ and $B_{\mathrm{AC}z}$) generate time-dependent components along $y$ and $z$, while a solenoid ($B_{\mathrm{AC}x}$) is used for the $x$ component. Several quadrupoles (not shown) help counteract static inhomogeneities.
@@
3 Methodology
The proposed methodology requires that the atomic magnetization adiabatically follows the magnetic field to be characterized, i.e., the atomic spins must precess much faster than changes in field orientation. With a time-dependent field that rotates at an angular speed $\omega$, the condition
$\gamma B\gg\omega$ (1)
must be fulfilled, where $\gamma$ is the gyromagnetic factor and $B$ is the modulus of the field. This requirement may not be compatible with the dressing experiments that inspired this work, for which $\omega\gg\gamma|B_{\mathrm{DC}}|$. Nevertheless, the calibration parameters determined with the proposed methodology can subsequently be applied also to operate under dressing conditions.
The calibration factors (valid for an assigned frequency $\omega/2\pi$) are determined by the accurate identification of coil currents (their relative phase and amplitudes) that must be applied to produce a circularly polarized time-dependent field that rotates at $\omega$. Specifically, we consider a field nominally made of a static component along $z$ and a rotating one on the $xy$ plane. Possible imperfections are considered, which may concern the orientation of the DC field and some degree of ellipticity in the polarization of the time-dependent field. Namely, we consider
$\vec{B}=B\hat{u}_{B}=B_{0}\big{(}\cos(\omega t)+m_{1},(1+\varepsilon)\sin(\omega t+\phi)+m_{2},m_{3}\big{)},$ (2)
with $m_{1},m_{2}$ accounting for spurious static field components on the polarization plane $xy$, and $\varepsilon,\phi$ for imperfections of the time-dependent field polarization. $B_{0}$ is the nominal amplitude of the rotating field and $m_{3}B_{0}$ is the nominal amplitude of the static one. The aim is to obtain a time-dependent field with circular polarization, i.e. $\varepsilon=\phi=0$, on a plane perpendicular to the static term, i.e. $m_{1}=m_{2}=0$.
In the hypothesis of adiabatic following, the macroscopic magnetization $\vec{M}$ maintains a constant modulus $M$ (apart from the decay due to the relaxation mechanisms) while precessing around $\vec{B}(t)$. Among the Cartesian components of $\vec{M}$, the one parallel to $\vec{B}$ remains approximately constant, while the perpendicular ones oscillate at the high frequency $\gamma|B|$ [26].
@@ 3 Methodology
The proposed procedure consists of a preliminary step aimed at ensuring $m_{1}=m_{2}=0$, and a main one to make $\varepsilon=\phi=0$. Both objectives are pursued through Fourier analysis of the Taylor approximation of $M_{\parallel x}(t)$ (eq. 3) – relevant details provided in the Appendix. The preliminary step pertains only to the DC components of the field and could alternatively be performed using traditional methods. The second step, which specifically addresses the AC components, leads to the primary scope of the proposed methodology.
3.1 Alignment of the static field along $z$ ($m_{1}=m_{2}=0$ condition)
A straightforward and commonly applied procedure to orient a static field along an axis (let it be $z$, as in our case) is based on minimizing the Larmor frequency by varying the currents that drive the transverse ($x$ and $y$) field components: the minimum is achieved when those transverse components of the field are fully compensated.
In the presence of a rotating field on the $xy$ plane, harmonic analysis of the detected signal, eq.A.1, can be used to reveal the presence of static components on the $xy$ plane, as well. Let’s analyze the harmonic content of the measured signal described by eq.3. The power ratio between the second-harmonic terms and the fundamental one (eqs.A.3 and A.2) is:
@@ 3.1 Alignment of the static field along $z$ ($m_{1}=m_{2}=0$ condition)
so that minimizing $A_{2}$ will lead to the condition $m_{1}=m_{2}=0$, i.e. to a static field perpendicular to the polarization plane of the time-dependent components. Similar to the case of Larmor frequency minimization, the transverse components appear quadratically in the quantity $A_{2}$ to be minimized.
@@
As shown in the Appendix, non-zero values of $\phi$ and $\varepsilon$ do not contribute to second harmonic terms, making the described procedure robust to imperfect polarization of the rotating field. Indeed (see eqs. A.3 and A.4), the polarization imperfections parametrized by $\phi$ and $\varepsilon$ produce only odd harmonic terms, as discussed in the next.
3.2 Refinement of the field circular polarization on the $xy$ plane ($\phi=\varepsilon=0$ condition)
Let’s assume that the static field (if any) has been aligned along $z$, i.e. let the condition $m_{1}=m_{2}=0$ and a generic $m_{3}$ be achieved. This paves the way to proceed with the second step of the proposed calibration method. Harmonics of the fundamental tone are now only ascribed to polarization imperfections, which, as shown in the Appendix, do contribute to even harmonics of the detected signal.
Making reference to eqs.A.2 and A.4, we may derive the power ratio between the third harmonic terms and the fundamental one, finding
@@ 3.2 Refinement of the field circular polarization on the $xy$ plane ($\phi=\varepsilon=0$ condition)
similarly to the case of eq.4, minimizing $A_{3}$ will lead toward the $\varepsilon=\phi=0$ condition, i.e. to a condition of perfectly circular polarization of the time-dependent field.
It is worth examining the possible consequences of imperfect perpendicularity of the static field on the accuracy of this procedure. In fact, besides the analyzed second-harmonic terms, a non-perpendicular static field would cause the emergence of third-harmonic term mimicking polarization issues. However, as shown in eq.A.4, in a first-order Taylor approximation $m_{1}$ and $m_{2}$ do not produce such an odd term, and they only contribute when also the second order (eq.A.5) is taken into account. In this sense, the procedure aimed to obtain circular polarization for the identification of $\varepsilon$ and $\phi$ is robust concerning the $m_{1}=m_{2}=0$ requirement.
4 Experimental validation
4.1 Static field alignment
We have measured a FPS after having aligned the field along $z$ by Larmor frequency minimization and in the absence of any rotating field. The static field inferred from the FPS is about 1 $\mu$T. We have then applied a field of 16 $\mu$T rotating at 3144 Hz, thus the nominal condition is $m_{1}=m_{2}=0,\,m_{3}=1/16$, the amplitudes and the relative phase of the oscillating $x,y$ components are set to $\pi/2$ based on the static calibration factors. The recorded signal is reported in Fig.4: just a weak second harmonic peak appears, and a third harmonic peak is visible, about 33 dB below the fundamental tone.